OpenStudy (anonymous):

Andrew estimated the weight of his dog to be 60 lb. The dog’s actual weight was 68 lb. What was the percent error in Andrew’s estimate? Round your answer to the nearest tenth of a percent. __________ %

OpenStudy (anonymous):

@ROBLOX

OpenStudy (amistre64):

what is 8 out of 60?

OpenStudy (amistre64):

or depending on how you read it 8 out of 68

OpenStudy (anonymous):

8.5?

OpenStudy (anonymous):

plz hurry

OpenStudy (firejay5):

@Sworewolf The correct answer is 11.7%, because \[\frac{ 8 }{ 68 }\] = 0.117 * 100 = 11.7%, but if rounded to next % it's 12% (I recommend round because it's accurate)